OpenStudy (ttp):

Which scenario has the system of linear equations: x plus y equals six…three x plus two y equals fourteen ? Katie bought six dollars and three dollars pastries at the bakery. She bought 5 pastries and spent fourteen dollars. Katie bought two dollars and six dollars pastries at the bakery. She bought 3 pastries and spent fourteen dollars. Katie bought three dollars and two dollars pastries at the bakery. She bought 5 pastries and spent fourteen dollars. Katie bought two dollars and three dollars pastries at the bakery. She bought 6 pastries and spent fourteen dollars.

jimthompson5910 (jim_thompson5910):

for me, it helps to write the equations out in numbers (instead of words), so let's do that first x+y = 6 3x+2y = 14

OpenStudy (ttp):

Okay x+y=6 3x+2y=14}?

OpenStudy (calculusxy):

D I believe since u are buying 6 pastries for $14.

jimthompson5910 (jim_thompson5910):

so if x was the number of 3 dollar pastries and y is the number of 2 dollar pastries, then x+y = 6 represents the idea that you are buying 6 pastries total

jimthompson5910 (jim_thompson5910):

if x = the number of 3 dollar pastries, then it will cost 3x dollars to buy x of them if y = the number of 2 dollar pastries, then it will cost 2y dollars to buy y of them so in total, it will cost 3x+2y dollars to buy x+y total pastries and because this total cost is 14, we say 3x+2y = 14

OpenStudy (ttp):

So D!

jimthompson5910 (jim_thompson5910):

calculusxy is correct